A 96 g of Potassium Bromide crystals is recovered by evaporating water from a solution with a
concentration of 45% at 60 degC. Assume that solution/mixture’s temperature at crystallization is
the same with the initial.
1-2) What is the initial molarity?
3-4) What is the initial concentration in w solute/100 w solvent?
5-6) Was the solution saturated? If not, what was it?
7-11) If the solution is 350 g initially, how many percent of the original solvent evaporated?
1
Expert's answer
2020-11-17T10:44:36-0500

45% means 45 g of salt in 100g solution.

And 450g of KBr in 1000 g solution. Since solvent is water density is 1g/cc


1-2) Molarity = (no. Of moles of KBr)/Volume

No. Of moles of KBr = 450/119 = 3.78moles

MOLARITY = 3.78moles/ (1L) = 3.78 molar


3-4)) mass of solute = 450g

Mass of solvent = 1000-450 g = 550g


In 100 g of solvent it will have 81.8 g of solute


5-6)It is not saturated solution. It is unsaturated , means more salt will be completely dissoolved.


7-11)Solution is 350 g initially out of which 96 g is recovered as KBr so remaining got evaporated it means 350-96 = 254 g solvent evaporated.

So (254×100/350) = 72.57% solvent got evaporated.